# Generation of Drug-Induced Cardiac Reactions towards Virtual Clinical Trials

### v1

@@ Abstract
Clinical trials are pivotal in cardiac drug development, yet they often fail due to inadequate efficacy and unexpected safety issues, leading to significant financial losses. Using in-silico trials to replace a part of physical clinical trials, e.g., leveraging advanced generative models to generate drug-influenced electrocardiograms (ECGs), seems an effective method to reduce financial risk and potential harm to trial participants. While existing generative models have demonstrated progress in ECG generation, they fall short in modeling drug reactions due to limited fidelity and inability to capture individualized drug response patterns. In this paper, we propose a Drug-Aware Diffusion Model (DADM), which could simulate individualized drug reactions while ensuring fidelity. To ensure fidelity, we construct a set of ordinary differential equations to provide external physical knowledge (EPK) of the realistic ECG morphology. The EPK is used to adaptively constrain the morphology of the generated ECGs through a dynamic cross-attention (DCA) mechanism. Furthermore, we propose an extension of ControlNet to incorporate demographic and drug data, simulating individual drug reactions. We compare DADM with the other eight state-of-the-art ECG generative models on two real-world databases covering $8$ types of drug regimens. The results demonstrate that DADM can more accurately simulate drug-induced changes in ECGs, improving the accuracy by at least $5.79\%$ and recall by $8\%$.
@@ 1 Introduction
In this paper, we design a Drug-Aware Diffusion Model (DADM), a novel deep-learning framework capable of generating post-dose ECGs based on pre-dose ECGs and clinical information. To generate realistic ECGs, we first construct a set of ODEs to describe EPK (McSharry et al., 2003). Then, we propose a dynamic cross-attention (DCA) mechanism for the denoising process, which can adaptively adjust the constraints imposed by the EPK according to the time steps. In the initial generation stage, the model should rely more on EPK to generate basic waveforms. In the later stage, the external control should be reduced to generate more detailed data. To accurately simulate the individual cardiac drug reactions on ECGs, we extend ControlNet (Zhang et al., 2023)to enable it to incorporate both EPK and tabular clinical data including demographic and drug information, which is named Clinical Information ControlNet (CICN).
We conduct experiments on two publicly available databases (Johannesen et al., 2014, 2016), comprising a total of 9,443 12-lead ECGs, which cover the effects of $8$ drug regimens to evaluate the performance of our method. Specifically, we evaluate the consistency of the generated ECGs and the real post-dose ECGs in terms of normality on three critical cardiac indicators. Compared to eight state-of-the-art (SOTA) ECG generative models, our proposed DADM achieves the best performance. Visualization results demonstrate that DADM can accurately simulate the effects of various drug reactions on ECGs across different subjects while maintaining fidelity.
@@ 3.2 Ordinary Differential Equation System
where $A=0.15\mathrm{mV}$. Baseline wander is a type of low-frequency noise ($0.05-3\mathrm{Hz}$ during stress testing) (Sörnmo and Laguna, 2005), caused by various factors: respiratory motion, patient movement, poor contact of electrode cables and ECG recording equipment, inadequate skin preparation at the electrode placement sites, and unclean electrodes (Golany et al., 2020). The solution of equation (1) is a trajectory in three-dimensional space with coordinates $(x(t),y(t),z(t))$. The generated ECG signal is only the third dimension, i.e., $z(t)$, which is the motion trajectory around a unit radius limit cycle on the $(x,y)$ plane. Each complete rotation of this circle corresponds to one heartbeat (or cardiac cycle). The locations of different points on the ECG signal (such as the $P$, $Q$, $R$, $S$, and $T$ waves) are determined by the parameters $\theta_{i}$, $a_{i}$, and $b_{i}$ in (1), where $i\in\{P,Q,R,S,T\}$. Parameters $\theta$, $a$ and $b$ determine fixed angles on the unit circle, amplitudes and durations, respectively. We denote the above 15 parameters by $\eta$ following (McSharry et al., 2003).
@@ 3.4 Dynamic Cross-Attention Mechanism
where $\mathrm{Attention_{i}}=\mathrm{softmax}(\mathbf{Q}_{i}\mathbf{K}^{T}/\sqrt{d})$, $i=\{1,2\}$. Here, $\mathbf{Q}_{1}=\mathbf{Q}_{ode}=\mathbf{W}^{q_{1}}\cdot\mathbf{y}_{ode}$, $\mathbf{Q}_{2}=\mathbf{Q}_{inter}=\mathbf{W}^{q_{2}}\cdot\mathbf{y}_{inter}$, $\mathbf{K}=\mathbf{W}^{k}\cdot\mathbf{y}_{ode}$, $\mathbf{V}=\mathbf{W}^{v}\cdot\mathbf{y}_{ode}$. And $\mathbf{W}^{q_{1}}$, $\mathbf{W}^{q_{2}}$, $\mathbf{W}^{k}$ and $\mathbf{W}^{v}$ are learnable projection matrices (Vaswani et al., 2017). $\alpha\in[0,1)$ is the result of a fully connected layer’s prediction for time-step $t$, used to measure the weight of EPK in the fusion process. The larger the value of $\alpha$, the greater the proportion of external EPK, and vice versa.
@@ 4.1 Dataset
We utilize two publicly available ECG datasets, including the ECGRDVQ database (Johannesen et al., 2014) and the ECGDMMLD database (Johannesen et al., 2016). These two datasets collectively comprise 9,443 12-lead ECGs from 44 subjects, covering the evaluation of the effects of $8$ types of drug regimens. The treatment regimens include: (1) Dofetilide; (2) Lidocaine + Dofetilide; (3) Mexiletine + Dofetilide; (4) Moxifloxacin + Diltiazem; (5) Placebo; (6) Quinidine; (7) Ranolazine; (8) Verapamil. We partition the data into training and testing sets with a $4:1$ ratio, ensuring that the proportion of ECGs affected by each drug in the training and testing sets also follows a $4:1$ ratio. More details can be found in the Technical Appendix.
@@ 4.2 Implementation Details
We train a generative model for each lead using $8$ NVIDIA GeForce RTX 4090 GPUs. The model training process is divided into two steps: (1) training a backbone DDPM which is injected EPK via DCA; (2) freezing the parameters of the DDPM and training the CICN. We set $T=1000$ and the forward process variances to constants increasing linearly from $\beta_{1}=10^{-4}$ to $\beta_{T}=0.02$. We use Adam optimizer (Diederik and Jimmy, 2015) with a batch size of $16$ and a learning rate of $2\times 10^{-4}$. We train all models for $50$ epochs. During the validation phase, we generate the corresponding post-dose ECGs at specific time points by varying the clinical data input to CICN. The code is uploaded as Supplementary Material.
@@ 4.3 Indicator Details
The calculation rules and normal ranges of each indicator are as follows: $\mathbf{QT_{c}}$ interval is the corrected $QT$ interval, representing the time from ventricular depolarization to complete repolarization, adjusted for heart rate. Here, we use the Bazett formula for correction: $QT_{c}=QT/\sqrt{RR}$, where $RR$ represents the heart rate. $QT_{c}\leq 450\mathrm{ms}$ in men and $QT_{c}\leq 470\mathrm{ms}$ in women are considered normal (Gupta et al., 2007). $\mathbf{PR}$ interval represents the time it takes for the electrical signal to travel from the atria to the ventricles, including atrial depolarization and the delay at the atrioventricular node. $120\mathrm{ms}\leq PR\leq 200\mathrm{ms}$ is normal (Kwok et al., 2016). $\mathbf{T_{peak}-T_{end}}$ interval represents the late phase of ventricular repolarization and reflects the dispersion of ventricular repolarization. $80\mathrm{ms}\leq T_{peak}-T_{end}\leq 113ms$ is considered normal (Icli et al., 2015).
@@ 4.5 Evaluation of Generated ECG Fidelity
The Fréchet Inception Distance (FID) (Heusel et al., 2017) is a traditional metric used for assessing the fidelity of generated images. Following Hu et al. (2024), we introduce the Fréchet ResNet Distance (FRD) as the metric to evaluate the generated ECG fidelity. Specifically, we utilize a 1D ResNet model (Hong et al., 2020) pre-trained on Challenge 2017 data (Clifford et al., 2017) instead of the Inception V3 (Szegedy et al., 2016) as the backbone for feature extraction from ECG data. We use the output from the network’s final fully connected layer as a feature vector to calculate the FRD on the test set. In addition, we adopt Root Mean Squared Error (RMSE) as a metric to evaluate the similarity between the generated and real ECGs post-dose. Lower FRD scores and RMSE indicate that the distribution of generated ECGs is closer to that of real ECGs.
We compare the average FRD scores and RMSE of 12-lead ECGs generated by different methods, and the results are presented in Table 3. Our findings demonstrate that our method achieves the optimal FRD score of $0.2431$ and an RMSE of $0.0127$.
Figure 4: (a) Ablation study with and without EPK, where w/o represents generating ECGs without EPK, Fixed represents fixed weight fusion, and Dyn represents DCA fusion; (b) The changing trend of $\alpha$ over time-steps.
@@ 4.6 Ablation Study
We conduct ablation studies to validate the importance of incorporating EPK and the effectiveness of DCA. We calculate the FRD scores of the generated ECGs under three conditions: without EPK, with fixed weight fusion, and with DCA. In the experiment with fixed weight fusion, we set $\alpha$ to be fixed at 0.5. For a fair comparison, the generative model with DCA does not have a CICN module. The experimental results are presented in Figure 4 (a), from which we have several observations: (1) The FRD achieved by DCA is significantly lower than that achieved by fixed weight fusion and fusion without EPK, demonstrating the effectiveness of our proposed DCA. (2) The FRD achieved by fixed weight fusion shows a slight improvement over fusion without EPK, indicating that merely introducing EPK does not substantially enhance the quality of ECG generation, which is consistent with the conclusion drawn in the third point of Table 1.
@@ 4.7 Visualization
To provide a more intuitive comparison of the generated ECGs with and without the introduction of EPK, we visualize the results, as shown in Figure 5. Note that in this experiment, we use the DCA mechanism to introduce EPK. Figure 5 shows that the model incorporating EPK not only captures the fundamental characteristics of heartbeats but also significantly reduces noise, thus enhancing the realism of the ECGs. In contrast, the model without EPK fails to synthesize clear waveforms, resulting in noisy signals. More visualization results can be found in the Technical Appendix.
@@
To more intuitively verify that DADM can accurately simulate the drug effects, we visualize ECGs showing the prolongation of the $PR$ interval caused by Verapamil’s effect on extending atrioventricular conduction time. Specifically, we visualize the ECGs of a subject one hour after administering Verapamil. The visualization results are presented in Figure 6. It is observed that after taking Verapamil, the subject’s $PR$ interval increases from $167\mathrm{ms}$ pre-dose to $293\mathrm{ms}$ post-dose. In the ECGs generated by our method, the $PR$ interval is $250\mathrm{ms}$. Both values fall within the abnormal range, demonstrating that our method can effectively simulate the reaction of drugs on ECGs. More visualization results can be found in the Technical Appendix.
@@ 5 Conclusions
In this paper, we design a novel framework named DADM to simulate cardiac drug reactions on ECGs for virtual clinical trials. To generate realistic ECGs, we first construct an ODE system to model EPK. Then, we propose a DCA mechanism that adaptively adjusts the constraints imposed by the EPK on ECG generation according to the time steps. To accurately simulate cardiac drug reactions, we propose a ControlNet extension called CICN to control the effects of different drugs on the ECGs according to the input clinical information. We conduct comparative and visualization experiments across two public datasets covering $8$ types of drug regimens, the results of which demonstrate superior performance over the other eight SOTA generative models in simulating cardiac drug reactions. The ablation study demonstrates the effectiveness of DCA mechanism.
@@ 6 Limitations
Appendix A Details of Datasets

### v2

@@ Abstract
Clinical trials remain critical in cardiac drug development but face high failure rates due to efficacy limitations and safety risks, incurring substantial costs. In-silico trial methodologies, particularly generative models simulating drug-induced electrocardiogram (ECG) alterations, offer a potential solution to mitigate these challenges. While existing models show progress in ECG synthesis, their constrained fidelity and inability to characterize individual-specific pharmacological response patterns fundamentally limit clinical translatability. To address these issues, we propose a novel Drug-Aware Diffusion Model (DADM). Specifically, we construct a set of ordinary differential equations to provide external physical knowledge (EPK) of the realistic ECG morphology. The EPK is used to adaptively constrain the morphology of the generated ECGs through a dynamic cross-attention (DCA) mechanism. Furthermore, we propose an extension of ControlNet to incorporate demographic and drug data, simulating individual drug reactions. Compared to the other eight state-of-the-art (SOTA) ECG generative models: 1) Quantitative and expert evaluation demonstrate that DADM generates ECGs with superior fidelity; 2) Comparative results on two real-world databases covering $8$ types of drug regimens verify that DADM can more accurately simulate drug-induced changes in ECGs, improving the accuracy by at least $5.79\%$ and recall by $8\%$. In addition, the ECGs generated by DADM can also enhance model performance in downstream drug-effect classification tasks.
@@ 1 Introduction
In this paper, we design a Drug-Aware Diffusion Model (DADM), a novel deep-learning framework capable of generating post-dose ECGs based on pre-dose ECGs and clinical information. To generate realistic ECGs, we first construct a set of ODEs to describe EPK (McSharry et al., 2003). Then, we propose a dynamic cross-attention (DCA) mechanism for the denoising process, which can adaptively adjust the constraints imposed by the EPK according to the time steps. In the initial generation stage, the model should rely more on EPK to generate basic waveforms. In the later stage, the external control should be reduced to generate more detailed data. To accurately simulate the individual cardiac drug reactions on ECGs, we extend ControlNet (Zhang et al., 2023)to enable it to incorporate both EPK and tabular clinical data, including demographic and drug information, which is named Clinical Information ControlNet (CICN).
We compare DADM with eight other SOTA ECG generative models on two publicly available databases (Johannesen et al., 2014, 2016), comprising a total of 9,443 12-lead ECGs, which cover the effects of $8$ drug regimens. The quantitative and expert assessment demonstrates that DADM generates ECGs with superior fidelity. The comparative results verify that DADM can more accurately simulate drug-induced changes in ECGs, improving the accuracy by at least $5.79\%$ and recall by $8\%$. Furthermore, the ECGs generated by DADM can also enhance model performance in downstream drug-effect classification tasks. And the ablation and visualization results demonstrate the effectiveness of each module.
@@ 3.2 Ordinary Differential Equation System
where $A=0.15\mathrm{mV}$. Baseline wander is a type of low-frequency noise ($0.05-3\mathrm{Hz}$ during stress testing) (Sörnmo and Laguna, 2005), caused by various factors: respiratory motion, patient movement, poor contact of electrode cables and ECG recording equipment, inadequate skin preparation at the electrode placement sites, and unclean electrodes (Golany et al., 2020). The solution of equation (1) is a trajectory in three-dimensional space with coordinates $(x(t),y(t),z(t))$. The generated ECG signal is only the third dimension, i.e., $z(t)$, which is the motion trajectory around a unit radius limit cycle on the $(x,y)$ plane. Each complete rotation of this circle corresponds to one heartbeat (or cardiac cycle). The locations of different points on the ECG signal (such as the $P$, $Q$, $R$, $S$, and $T$ waves) are determined by the parameters $\theta_{i}$, $a_{i}$, and $b_{i}$ in (1), where $i\in\{P,Q,R,S,T\}$. Parameters $\theta$, $a$ and $b$ determine fixed angles on the unit circle, amplitudes and durations, respectively. We kindly refer the readers to Appendix A for more details about the parameters.
@@ 3.4 Dynamic Cross-Attention Mechanism
where $\mathrm{Attention_{i}}=\mathrm{softmax}(\mathbf{Q}_{i}\mathbf{K}^{T}/\sqrt{d})$, $i=\{1,2\}$. Here, $\mathbf{Q}_{1}=\mathbf{Q}_{ode}=\mathbf{W}^{q_{1}}\cdot\mathbf{y}_{ode}$, $\mathbf{Q}_{2}=\mathbf{Q}_{inter}=\mathbf{W}^{q_{2}}\cdot\mathbf{y}_{inter}$, $\mathbf{K}=\mathbf{W}^{k}\cdot\mathbf{y}_{ode}$, $\mathbf{V}=\mathbf{W}^{v}\cdot\mathbf{y}_{ode}$. And $\mathbf{W}^{q_{1}}$, $\mathbf{W}^{q_{2}}$, $\mathbf{W}^{k}$ and $\mathbf{W}^{v}$ are learnable projection matrices (Vaswani et al., 2017). $\alpha\in[0,1)$ is the result of a fully connected layer’s prediction for time-step $t$, used to measure the weight of EPK in the fusion process. The larger the value of $\alpha$, the greater the proportion of external EPK, and vice versa. Although the design of DCA is straightforward, its innovation lies in dynamically adjusting the constraints of EPK without introducing redundant operations, which is well-suited for our task. The comparative experiments with other attention mechanisms are presented in Section 4.7.
@@ 4.1 Dataset
We utilize two publicly available ECG datasets, including the ECGRDVQ database (Johannesen et al., 2014) and the ECGDMMLD database (Johannesen et al., 2016). These two datasets collectively comprise 9,443 12-lead ECGs from 44 subjects, covering the evaluation of the effects of $8$ types of drug regimens. The treatment regimens include: (1) Dofetilide; (2) Lidocaine + Dofetilide; (3) Mexiletine + Dofetilide; (4) Moxifloxacin + Diltiazem; (5) Placebo; (6) Quinidine; (7) Ranolazine; (8) Verapamil. We partition the data into training and testing sets with a $4:1$ ratio, ensuring that the proportion of ECGs affected by each drug in the training and testing sets also follows a $4:1$ ratio. More details can be found in Appendix B.
@@ 4.2 Implementation Details
We train a generative model for each lead using $8$ NVIDIA RTX 4090 GPUs. The training process is divided into two steps: (1) train a backbone DDPM that is injected with EPK via DCA; (2) train the CICN while freezing the parameters of DDPM. We set $T=1000$ and the forward process variances to constants increasing linearly from $\beta_{1}=10^{-4}$ to $\beta_{T}=0.02$. We use Adam optimizer (Diederik and Jimmy, 2015) with a batch size of $16$ and a learning rate of $2\times 10^{-4}$. We train all models for $50$ epochs. During the validation phase, we generate the corresponding post-dose ECGs at specific time points by varying the clinical data input to CICN.
@@ 4.3 Indicator Details
The calculation rules and normal ranges of each indicator are as follows: $\mathbf{QT_{c}}$ interval is the corrected $QT$ interval, representing the time from ventricular depolarization to complete repolarization, adjusted for heart rate. Here, we use the Bazett formula for correction: $QT_{c}=QT/\sqrt{RR}$, where $RR$ represents the heart rate. $QT_{c}\leq 450\mathrm{ms}$ in men and $QT_{c}\leq 470\mathrm{ms}$ in women are considered normal (Gupta et al., 2007). $\mathbf{PR}$ interval represents the time it takes for the electrical signal to travel from the atria to the ventricles, including atrial depolarization and the delay at the atrioventricular node. $120\mathrm{ms}\leq PR\leq 200\mathrm{ms}$ is normal (Kwok et al., 2016). $\mathbf{T_{peak}-T_{end}}$ interval represents the late phase of ventricular repolarization and reflects the dispersion of ventricular repolarization. $80\mathrm{ms}\leq T_{peak}-T_{end}\leq 113ms$ is considered normal (Icli et al., 2015). The onset and offset points of the three indicators are determined by ten cardiologists with at least five years of experience.
@@ 4.5 Evaluation of Generated ECG Fidelity
Quantitative Evaluation.
The Fréchet Inception Distance (FID) (Heusel et al., 2017) is a traditional metric used for assessing the fidelity of generated images. Following Hu et al. (2024), we introduce the Fréchet ResNet Distance (FRD) as the metric to evaluate the generated ECG fidelity. Specifically, we utilize a 1D ResNet model (Hong et al., 2020) pre-trained on Challenge 2017 data (Clifford et al., 2017) instead of the Inception V3 (Szegedy et al., 2016) as the backbone for feature extraction from ECG data. We use the output from the network’s final fully connected layer as a feature vector to calculate the FRD on the test set. In addition, we adopt Root Mean Squared Error (RMSE) as a metric to evaluate the similarity between the generated and real ECGs post-dose. Lower FRD scores and RMSE indicate that the distribution of generated ECGs is closer to that of real ECGs. We compare the average FRD scores and RMSE of 12-lead ECGs generated by different methods, and the results are presented in Table 3. Our findings demonstrate that our method achieves the optimal FRD score of $0.2431$ and an RMSE of $0.0127$.
Expert Evaluation.
We select $100$ generated ECGs and $100$ real ECGs of lead II, and then invite ten cardiologists with at least five years of experience to distinguish whether they are real or generated. The results show that $93.80\%\pm 2.66\%$ (mean±$\pm$±std) of the generated ECGs are considered to be real, demonstrating the realism of our generated ECGs.
Figure 4: (a) Ablation study of attention mechanisms; (b) Changing trend of $\alpha$ over time-steps.
4.6 Evaluation of Downstream Classification Task
To validate that the ECGs generated by our method improve the performance of classification models in downstream tasks, we design a binary classification experiment for drug effects (specifically Quinidine). We train four ResNet-1D models using the following datasets to classify whether the drug is administered: 1) $1,000$ real ECGs, 2) $1,000$ real $+$ $100$ generated ECGs, 3) $1,000$ real $+$ $200$ generated ECGs, and 4) $1,000$ real $+$ $500$ generated ECGs. The test set is identical for all models, consisting of $100$ pre-dose and $100$ post-dose ECGs. The results are shown in Table 4, demonstrating that our generated data helps improve the performance of drug classification models. More importantly, the cost of generating data is significantly lower than that of acquiring real data. Our method provides a viable solution to the problem of data scarcity in real medical scenarios.
Dataset Accuracy
$1,000$ real ECGs $92.00\%$
$+100$ generated ECGs $94.00\%$
$+200$ generated ECGs $95.50\%$
$+500$ generated ECGs $97.75\%$
Table 4: Evaluation of Quinidine efficacy classification accuracy with varying numbers of generated ECGs.
@@ 4.7 Ablation Study
To validate the importance of EPK and the effectiveness of DCA, we compare the FRD scores of ECGs generated using DCA with those generated using two baseline attention mechanisms and two widely-used attention mechanisms in time series analysis, including no attention (NA), fixed-weight attention (FWA), prob-sparse attention (PSA) (Zhou et al., 2021), and exogenous-to-endogenous cross-attention (EECA) (Wang et al., 2024). NA means without fusing EPK, and the $\alpha$ of FWA is fixed at $0.5$. For a fair comparison, the generative model with DCA does not have a CICN module. The experimental results are presented in Figure 4 (a), from which we have several observations: (1) Generated ECGs integrated with EPK exhibit lower FRD scores compared to those without EPK fusion, proving the effectiveness of incorporating EPK. (2) DCA outperforms other static fusion mechanisms, demonstrating that dynamically adjusting EPK’s constraints is effective and particularly well-suited for our task.
@@ 4.8 Visualization
To provide a more intuitive comparison of the generated ECGs with and without the introduction of EPK, we visualize the results, as shown in Figure 5. Note that in this experiment, we use the DCA mechanism to introduce EPK. Figure 5 shows that the model incorporating EPK not only captures the fundamental characteristics of heartbeats but also significantly reduces noise, thus enhancing the realism of the ECGs. In contrast, the model without EPK fails to synthesize clear waveforms, resulting in noisy signals. More visualization results can be found in Appendix C.
@@
To more intuitively verify that DADM can accurately simulate the drug effects, we visualize ECGs showing the prolongation of the $PR$ interval caused by Verapamil’s effect on extending atrioventricular conduction time. Specifically, we visualize the ECGs of a subject one hour after administering Verapamil. The visualization results are presented in Figure 6. It is observed that after taking Verapamil, the subject’s $PR$ interval increases from $167\mathrm{ms}$ pre-dose to $293\mathrm{ms}$ post-dose. In the ECGs generated by our method, the $PR$ interval is $270\mathrm{ms}$. Both values fall within the abnormal range, demonstrating that our method can effectively simulate the reaction of drugs on ECGs. More visualization results can be found in Appendix C.
@@ 5 Conclusions
In this paper, we design a novel framework named DADM to simulate cardiac drug reactions on ECGs for virtual clinical trials. To generate realistic ECGs, we first construct an ODE system to model EPK. Then, we propose a DCA mechanism that adaptively adjusts the constraints imposed by the EPK on ECG generation according to the time steps. To accurately simulate cardiac drug reactions, we propose a ControlNet extension called CICN to control the effects of different drugs on the ECGs according to the input clinical information. We conduct experiments across two public datasets covering $8$ types of drug regimens, the results of which demonstrate superior performance of DADM over the other eight SOTA generative models in ensuring fidelity and simulating drug reactions. Furthermore, the ECGs generated by DADM can also enhance model performance in downstream drug-effect classification tasks. And the ablation and visualization results demonstrate the effectiveness of each module.
@@ 6 Limitations
Acknowledgements
We extend our sincere gratitude to Dr. Zheng Xiangshang and other cardiologists for their timely assistance in ECG signal analysis, including the determination of interval onset and offset locations, and the authenticity assessment of generated ECG signals.
Appendix A Details of ODE Parameters
The ODE system is defined by $3$ parameter groups, including $\theta_{i}$: Phase positions governing wave event localization; $a_{i}$: Amplitude parameters controlling the wave morphology; $b_{i}$: Temporal scaling factors regulating the wave duration, where $i\in\{P,Q,R,S,T\}$. To concretely demonstrate this parametric control, Figure 7 visualizes a prototypical normal heartbeat with clinically validated reference values: $(\theta_{P},\theta_{Q},\theta_{R},\theta_{S},\theta_{T})=(-\frac{1}{3}\pi,-\frac{1}{12}\pi,0,\frac{1}{12}\pi,\frac{1}{2}\pi)$, $(a_{P},a_{Q},a_{R},a_{S},a_{T})=(1.2,-5.0,30.0,-7.5,0.75)$, and $(b_{P},b_{Q},b_{R},b_{S},b_{T})=(0.25,0.1,0.1,0.1,0.4)$. We assign values to these parameters following McSharry et al. [2003].
Figure 7: Visualization of the heartbeat in the 3-D space. The dashed line reflects the limit cycle of unit radius while the black points show the positions of the $P$, $Q$, $R$, $S$, and $T$ wave events.
Appendix B Details of Datasets